The final answer should be in the format of "tolerance of ± in the third significant digit."


A car is traveling around a circular track of a 730-ft radius. If the magnitude of its total acceleration is 11 ft/sec2 at the instant when its speed is 50 mi/hr, determine the rate at which the car is changing its speed.

at = ± _________ ft/sec2


Write the vector expression for the acceleration a of the mass center G of the simple pendulum in both n-t and x-y coordinates for the instant when θ = 33° if θ˙ = 2.41 rad/sec and θ¨ = 4.445 rad/sec2.

a = (_____ en + _____ et) ft/sec2

a = (_____ i + ______ j) ft/sec2





1
Expert's answer
2021-04-01T18:41:38-0400

1)

"a_t=a-\frac{v^2}{r}=10.997~\frac{ft}{s^2},"

2)

"a_n=\frac{v^2}{r}=r\theta'^2=v\theta',"

"a_t=v'=r\theta'',"

"a_n=1.467\cdot 2.41=3.535e_n,"

"a_t=730\cdot 4.445=3244.850e_t,"

"i=-a_tsin\theta-a_ncos\theta=-1770.237,"

"j=a_t cos\theta-a_n sin \theta=2719.435."